Y=X^1/5Z^4/5.what is expansion path equation?
1
Expert's answer
2020-02-07T10:02:52-0500

If Y=X^1/5Z^4/5, then expansion path equation is:

"\frac{MPx}{MPz} = \frac{Px}{Pz},"

"MPx = Y'(x) = 0.2(Z\/X)^{4\/5} ,"

"MPz = Y'(z) = 0.8(X\/Z)^{1\/5},"

"\frac{0.2(Z/X)^{4/5}}{0.8(X/Z)^{1/5}} = \frac{Px}{Pz},"

"0.25(\frac{Z}{X})^{4/5}= \frac{Px}{Pz}."